OpenStudy (anonymous):

here it is: Which system of equations can be used to solve the following problem? Each child ticket for a ride costs $2, while each adult ticket costs $6. If the ride collected a total of $148, and 38 tickets were sold, how many of each type of ticket were sold? Let c be the number of child tickets and a be the number of adult tickets.

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

A. 2c+6a=148 a+c=38 B. 2a+6c=38 a+c=148 C. 6c+2a=148 c+a=38 D. 2c+6a=38 a+c=148

OpenStudy (anonymous):

those are my options

OpenStudy (catlover5925):

first EQUATION --->2c + 6a = 148 C=child and A= adult already knowing that you know the answer is A because child tickets were $2 and child was c, so it cant be B because it says 2a...and it cant be D because it put # of tickets at the end when it has to be money FINAL ANSWER ----> A

OpenStudy (anonymous):

wow, you are amazing at this :D and a good explainer

OpenStudy (catlover5925):

thanks, tag me anytime if u need help

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@cutegirl